[figure: 9.7 cm]

NEUTRON DIFFRACTION STUDY OF SEMI-HEUSLER SOLID

SOLUTIONS.

K. Kaczmarska ^{a}, J. Pierre^{b}, B. Ouladdiaf^{c}

a Institute of Physics, University of Silesia, Uniwersytecka 4, 40007 Katowice, Poland

b Laboratoire Louis Néel, CNRS, 166X, 38042 Grenoble, France c Institute Laue Langevin, 156X, 38042 Grenoble, France

Recent studies of CoV_{1−x}Mn_{x}Sb and NiTi_{1−x}Mn_{x}Sb solid solutions show very interesting properties. For low Mn concentration, 3d states of manganese are well localized, the resistivity is metallic. Magnetic interactions strongly increase near x = 0.4, while semiconducting-like behaviour is observed in the range x = 0.35-0.45. These features are explained by the Anderson localization of d states in these disordered solutions. In order to investigate the atomic and magnetic structures and disorder, neutron diffraction experiments have been done. The crystallographic order is discussed together with the X-ray data. Ferromagnetic order followed by short range order has been observed for CoV_{1−x}Mn_{x}Sb samples with x from 0.25 to 0.45. Spin-glass behaviour has been found for NiTi_{1−x}Mn_{x}Sb samples with x = 0.35, 0.45 with spin-glass temperatures below 15K. The phase diagram of these solutions is discussed.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : oral

Corresponding author : K.Kaczmarska

Address for correspondence : Instytut Fizyki U.Sl. Uniwersytecka 4 40-007 Katowice

Email address : kaczmars@us.edu.pl